You want to purchase a $200,000 house and have a 20% down payment. You enter into a 30 year mortgage at a rate of 4.5%. What is your monthly payment?

Show inputs in financial calculator
Answer #1

The amount of mortgage after paying the down payment is = $160,000 ($2,00,000*0.8)

So, PV = $160,000

N = 30*12 = 360 (since 1 year has 12 months, 30 years has 360 monthly payments )

I/Y = 0.375% (4.5%/12)

So, compute the PMT = ($810.6965)

So, the monthly payments are 810.6965